# Excited State Densities from Time-Dependent Density Functional Response Theory

### v1

@@ 1 Introduction
The advent of time-dependent density functional theory (TDDFT) forty years ago 1, 2, 3 has enabled the calculation of electronic spectra, response properties, and general non-perturbative dynamics, for large and complex systems that would otherwise not be possible. This is due to the one-to-one mapping between time-dependent potentials and densities, for a specified initial state, opening the possibility of finding a non-interacting (Kohn-Sham) system that reproduces the same density while evolving in the computationally simpler one-body Schrödinger equation. While the approach would formally yield exact observables of the interacting system, in practice approximations must be made for the exchange-correlation potential. The majority of applications operate in the linear response regime, where one considers a time-dependent perturbation on the system, from whose response excitation energies and transition properties out of the ground-state can be extracted, and, with the approximations in use today, an unprecedented balance between accuracy and efficiency has been achieved. When thinking about the linear response of a ground-state system, it can be understood that properties related to transition densities between the ground and excited states of the true interacting system, $\langle\Psi_{I}|\hat{n}({\bf r})|\Psi_{0}\rangle$, can be directly obtained, but it is not immediately obvious that excited-state densities $\langle\Psi_{I}|\hat{n}({\bf r})|\Psi_{I}\rangle$ can be extracted from the formalism. However, with the recognition that excited states are stationary points of the Hamiltonian, then using time-independent perturbation theory instead of time-dependent perturbation theory offers a route to extracting their densities from energies computed from linear response; equivalently, framing the problem as a variational one 4, 5. (We note that excited state densities can also be obtained from extrema of the purely the ground-state energy functional 6, 7).
@@
In this paper, we work directly in real space instead of in the space of single-excitations that the earlier pioneering work was framed in 4, 5, and going beyond the use of an adiabatic kernel that the earlier work was limited to. The latter extension enables us to obtain densities for states of double-excitation character. We explore the small-matrix approximation limit of our expressions on a series of one-dimensional model systems that cover local excitations, charge-transfer excitations, and double excitations. Motivated by the works8, 9, 10, 11, 12, 13, 14, 15 suggesting that the hybrid exchange-correlation functionals with a large fraction of exact exchange yield accurate excited-state properties, we explore the performance of the excited-state densities in two opposite limits of approximation – in the adiabatic local density approximation (LDA), and adiabatic exact exchange (EXX). We find that the choice of ground-state approximation is important in getting good Kohn-Sham densities in the first place, on top of which TDDFT provides corrections that bring the density close to the exact. These corrections involve a sum over all orbitals, whose accuracy depends on the choice of exchange-correlation functional, and we study the convergence of this sum. Finally, we show that the DSMA and DSPA frequency-dependent approximations provide good approximations to the densities of states of double-excitation character, while the adiabatic approximation fails.
@@ 2 Excited-state densities from time-independent perturbation theory
A standard result from time-independent perturbation theory of quantum mechanics is that the first-order change to an excited-state energy is the expectation value of the perturbation evaluated in that excited state. Applied to many-electron systems when this perturbation is in the externally applied one-body potential, this then yields $\delta E_{I}=\int n_{I}({\bf r})\delta v_{\rm ext}({\bf r})d^{3}r$, where $n_{I}({\bf r})=\sum_{\sigma_{1}..\sigma_{N}}\int|\Psi_{I}({\bf r}\sigma_{1},{\bf r}_{2}\sigma_{2}...{\bf r}_{N}\sigma_{N})|^{2}d^{3}r_{2}..d^{3}r_{N}$ is the density of the unperturbed excited state $I$. Thus, knowledge of the excited state energy and how it varies with the potential yields the excited-state density itself:
@@
Eq. 1 provides the cornerstone for obtaining excited state densities from TDDFT response: while ground-state DFT provides the ground-state energy $E_{0}$, the TDDFT linear response formalism provides excitation energies $\omega_{I}$, so taking the functional derivative with respect to the external potential of the sum of these yields the excited-state density, exact in principle and dependent on functional approximations made in practise, as we shall see shortly. Thus, we write the excited state energy in two parts:
@@
[…]
@@
where the matrix elements of $\Omega(\omega)$ are expressed in the single-excitation basis, $q=i\rightarrow a,\;q^{\prime}=j\rightarrow b$, as
@@
and $\Phi_{q}(\mathbf{x})=\phi_{i}(\mathbf{x})\phi_{a}(\mathbf{x})$ represents a KS single-excitation transition density.
Evaluating the derivative in Eq. (1) from Eq. 2 we have
@@
obtaining the derivative from Eq. (3) in the following way. Multiplying by $\mathbf{G}_{I}^{\dagger}$ on the left, and then taking the derivative, yields the left-hand side as
@@
$\Delta n_{I}({\bf r})=\frac{1}{2\omega_{I}}\mathbf{G}_{I}^{\dagger}\frac{\delta\Omega(\omega_{I})}{\delta v_{\rm ext}({\bf r})}\left(\mathbf{G}_{I}^{\dagger}\right)^{-1}=\frac{1}{2\omega_{I}}\mathbf{G}_{I}^{-1}\frac{\delta\Omega(\omega_{I})}{\delta v_{\rm ext}({\bf r})}\mathbf{G}_{I}$ (11)
This expression gives us the exact density-difference, once we compute the functional derivative of the TDDFT matrix (see shortly).
While Eq. (11) is independent of the choice of normalization of the response eigenvectors $\mathbf{G}_{I}$ (as it must be from Eq. (8)), it may be useful to consider the particular normalization derived in Ref. 16 which directly relates the eigenvectors to the oscillator strengths via $f_{I}=\frac{2}{3}\omega_{I}|\langle\Psi_{0}|{\bf r}|\Psi_{I}\rangle|^{2}=\frac{2}{3}\left|\sum_{q,q^{\prime}}\mathbf{r}_{q}^{\dagger}S_{q,q^{\prime}}^{-1/2}G_{I,q^{\prime}}\right|^{2}$ 17, where $S$ is the diagonal matrix $S_{qq^{\prime}}=\delta_{qq^{\prime}}\frac{1}{v_{q}}$. Ref. 16 showed that the required normalization condition is
@@
With this choice, we have
$\Delta n_{I}({\bf r})=\frac{1}{2\omega_{I}}\frac{\delta\omega^{2}_{I}}{\delta v_{\text{ext}}({\bf r})}=\frac{1}{2\omega_{I}}\left(\mathbbm{1}+\mathbf{G}_{I}^{\dagger}\left.\frac{\partial\Omega(\omega)}{\partial(\omega^{2})}\right|_{\omega=\omega_{I}}\mathbf{G}_{I}\right)^{-1}\mathbf{G}_{I}^{\dagger}\frac{\delta\Omega(\omega_{I})}{\delta v_{\text{ext}}({\bf r})}\mathbf{G}_{I}.$ (13)
We note that the functional derivative $\frac{\delta\Omega(\omega_{I})}{\delta v_{\rm ext}({\bf r})}$ has two contributions evident from Eq. (4): one is from the variation of KS orbitals and eigenvalues $\{\phi_{i}({\bf r}),\epsilon_{i}\}$ that explicitly appear in Eq. (4) and from the variation of the functional-argument $n_{0}$ of the $f_{\scriptscriptstyle\rm XC}$ kernel under small changes of the external potential, and the other is from the possible frequency-dependence of the xc kernel $f_{\scriptscriptstyle\rm XC}(\omega)$ because this is evaluated at the solution point of Eq. (3), $\omega_{I}$, which depends on $v_{\rm ext}({\bf r})$, $\frac{\delta f_{\scriptscriptstyle\rm XC}(\omega_{I})}{\delta v_{\rm ext}({\bf r})}=\left.\frac{\delta f_{\scriptscriptstyle\rm XC}(\omega)}{\delta\omega}\right|_{\omega=\omega_{I}}\frac{\delta\omega_{I}}{\delta v_{\rm ext}({\bf r})}$. Within the commonly-used adiabatic approximation, the TDDFT matrix is frequency-independent and so $\frac{\delta f_{\scriptscriptstyle\rm XC}(\omega)}{\delta\omega}=0$ and $\frac{\partial\Omega}{\partial(\omega^{2})}$ in Eq. 13 is zero. In the general case, frequency-dependence of the exact kernel brings non-linearity to Eq. (3), resulting therefore in more solutions than the size of the KS single-excitation matrix. This is essential to correctly get states corresponding to states of double-excitation character 20, 21, 22, 23, 24, which are not accessible with an adiabatic approximation. While the frequency-dependence of the $f_{\scriptscriptstyle\rm XC}$ kernel also affects the density-difference dependence through Eq. (13), this equation is, however, still linear, as the matrix is evaluated at the solution frequency $\omega_{I}$, yielding a unique and well-defined density-difference in Eq. 8.
A simplification to Eq. (8) arises once we separate the two contributions to the functional derivative of the $f_{\scriptscriptstyle\rm XC}$ kernel that were mentioned above:
@@
where the second term is only non-zero in the case of non-adiabatic functionals. Transferring this separation to the full $\Omega(\omega)$ matrix, and inserting into Eq. (8), we find
@@
We note that this does not trivially follow from Eq. 11 because in general $G_{I}$ need not be chosen orthonormal; it certainly is not in the case of the normalization chosen for oscillator strengths, Eqs..
Eq. 11, or Eq. 15 define the exact excited-state densities once the terms of the derivative matrix $\frac{\delta\Omega_{qq^{\prime}}(\omega)}{\delta v_{\rm ext}({\bf r})}$ are calculated. When KS states are considered real, one finds:
@@
$+4\sqrt{\nu_{q}\nu_{q^{\prime}}}\left(\sum_{p\neq i}^{\infty}\frac{1}{\epsilon_{i}-\epsilon_{p}}f_{\mathrm{HXC},\,pa,\,jb}\left(\omega\right)\Phi_{ip}(\mathbf{x})+\sum_{p\neq a}^{\infty}\frac{1}{\epsilon_{a}-\epsilon_{p}}f_{\mathrm{HXC},\,ip,\,jb}\left(\omega\right)\Phi_{pa}(\mathbf{x})\right.$
$\left.\left.+\sum_{p\neq j}^{\infty}\frac{1}{\epsilon_{j}-\epsilon_{p}}f_{\mathrm{HXC},\,ia,\,pb}\left(\omega\right)\Phi_{jp}(\mathbf{x})+\sum_{p\neq b}^{\infty}\frac{1}{\epsilon_{b}-\epsilon_{p}}f_{\mathrm{HXC},\,ia,\,jp}\left(\omega\right)\Phi_{pb}(\mathbf{x})\right)\right\}\left(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}\right)^{-1}(\mathbf{x},{\bf r})$
@@
Section 1 in the Supplementary Material contains the details of the derivation of Eq. 16. Here we have used the following short-hands. First, we note that when $\chi,\chi_{\scriptscriptstyle\rm S},f_{\scriptscriptstyle\rm HXC}$ are written without any frequency-dependence, this indicates they are to be evaluated at the static limit $\omega=0$: $\chi({\bf r},{\bf r}^{\prime})=\frac{\delta n({\bf r})}{\delta v_{\rm ext}({\bf r}^{\prime})}|_{n=n_{0}},\chi_{\scriptscriptstyle\rm S}({\bf r},{\bf r}^{\prime})=\frac{\delta n({\bf r})}{\delta v_{\scriptscriptstyle\rm S}({\bf r}^{\prime})}|_{n=n_{0}},f_{\scriptscriptstyle\rm HXC}=f_{\scriptscriptstyle\rm HXC}[n_{0}]({\bf r},{\bf r}^{\prime})=\frac{\delta v_{\scriptscriptstyle\rm HXC}[n]({\bf r})}{\delta n({\bf r}^{\prime})}|_{n=n_{0}}$. In contrast, $f_{\scriptscriptstyle\rm HXC}(\omega)=f_{\scriptscriptstyle\rm HXC}[n_{0}]({\bf r},{\bf r}^{\prime},\omega)$ is defined as in Eq. (6). The KS approximation to the single-excitation density difference is denoted $\Delta n^{\scriptscriptstyle\rm KS}_{q}({\bf x})=\phi_{a}^{2}({\bf x})-\phi_{i}^{2}({\bf x})$, while $\Phi_{mn}({\bf x})=\phi_{m}({\bf x})\phi_{n}({\bf x})$ is defined in the same way as the the KS transition density earlier was, except now generalized since $\phi_{m},\phi_{n}$ could both be occupied or unoccupied orbitals. Note that $p$ in the sums run over all states, occupied and unoccupied, so this expression probes the xc kernel matrix elements (Eq. 5) in a much larger range than needed in the calculation of the excitation spectra; the latter involves matrix elements only between single excitations. The real-space matrix $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}({\bf x},{\bf r})$ is defined using the static $f_{\scriptscriptstyle\rm HXC}$ kernel and KS response function, and the product $f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}$ indicates the convolution, $f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}({\bf x},{\bf r})=\int d^{3}x^{\prime}f_{\scriptscriptstyle\rm HXC}({\bf x},{\bf x}^{\prime})\chi_{\scriptscriptstyle\rm S}({\bf x}^{\prime},{\bf r})$. Finally, $\tilde{g}_{{\scriptscriptstyle\rm HXC},qq^{\prime}}({\bf x})=\int d^{3}x^{\prime}d^{3}x^{\prime\prime}\tilde{g}_{\scriptscriptstyle\rm XC}[n_{0}](\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime},\mathbf{x},\omega)\Phi_{q}({\bf x}^{\prime})\Phi_{q^{\prime}}({\bf x}^{\prime\prime})$ is the matrix in KS single-excitation space of
@@
Eqs. 1516, once added to the ground-state density, give in-principle the exact excited-state density for the state $I$. As such Eq.(16) must integrate to zero. To see that it does, consider the Taylor expansions of the inverse
@@
We note that Refs. 25, 26, 4, 5 begin instead with finding stationary points of a Lagrangian that expresses the variational property of the excited state energy, working in the KS single-excitation subspace to directly obtain matrix equations for the so-called $Z$-vector that contains relaxation contributions to a difference density matrix; solving these equations yields excited state gradients and other properties. Our Eqs. 1516 are equivalent but instead yields the densities directly in real-space, and moreover, applies also to the case of non-adiabatic kernels.

### v2

@@ 1 Introduction
The advent of time-dependent density functional theory (TDDFT) forty years ago 1, 2, 3 has enabled the calculation of electronic spectra, response properties, and general non-perturbative dynamics, for large and complex systems that would otherwise not be possible. This is due to the one-to-one mapping between time-dependent potentials and densities, for a specified initial state, opening the possibility of finding a non-interacting (Kohn-Sham) system that reproduces the same density while evolving in the computationally simpler one-body Schrödinger equation. While the approach would formally yield exact observables of the interacting system, in practice approximations must be made for the exchange-correlation potential. The majority of applications operate in the linear response regime, where one considers a time-dependent perturbation on the system, from whose response excitation energies and transition properties out of the ground-state can be extracted, and, with the approximations in use today, an unprecedented balance between accuracy and efficiency has been achieved. When thinking about the linear response of a ground-state system, it can be understood that properties related to transition densities between the ground $\Psi_{0}$ and excited states $\Psi_{I}$ of the true interacting system, $\langle\Psi_{I}|\hat{n}({\bf r})|\Psi_{0}\rangle$, can be directly obtained, $\hat{n}({\bf r})$ being the density operator, but it is not immediately obvious that excited-state densities $\langle\Psi_{I}|\hat{n}({\bf r})|\Psi_{I}\rangle$ can be extracted from the formalism. However, with the recognition that excited states are stationary points of the Hamiltonian, then using time-independent perturbation theory instead of time-dependent perturbation theory offers a route to extracting their densities from energies computed from linear response; equivalently, framing the problem as a variational one 4, 5. (We note that excited state densities can also be obtained from extrema of the purely the ground-state energy functional 6, 7, generalized adiabatic connection formalism 8, 9, stationary principles based on a bi-functional of the excited and ground state density 10, 11, and ensemble DFT approaches 12, 13, 14, 15).
@@
In this paper, we work directly in real space instead of in the space of single-excitations that the earlier pioneering work was framed in 4, 5, and going beyond the use of an adiabatic kernel that the earlier work was limited to. The latter extension enables us to obtain densities for states of double-excitation character. We explore the small-matrix approximation limit of our expressions on a series of one-dimensional model systems that cover local excitations, charge-transfer excitations, and double excitations. Motivated by the works16, 17, 18, 19, 20, 21, 22, 23 suggesting that the hybrid exchange-correlation functionals with a large fraction of exact exchange yield accurate excited-state properties, we explore the performance of the excited-state densities in two opposite limits of approximation – in the adiabatic local density approximation (LDA), and adiabatic exact exchange (EXX). We find that the choice of ground-state approximation is important in getting good Kohn-Sham densities in the first place, on top of which TDDFT provides corrections that bring the density close to the exact. These corrections involve a sum over all orbitals, whose accuracy depends on the choice of exchange-correlation functional, and we study the convergence of this sum. Finally, we show that the dressed small matrix (DSMA) and dressed single pole (DSPA) approximations for frequency-dependent kernel provide good approximations to the densities of states of double-excitation character, while the adiabatic approximation fails.
@@ 2 Excited-state densities from time-independent perturbation theory
A standard result from time-independent perturbation theory of quantum mechanics is that the first-order change to an excited-state energy $\delta E_{I}$} is the expectation value of the perturbation evaluated in that excited state. Applied to many-electron systems when this perturbation is in the externally applied one-body potential $v_{\rm ext}({\bf r})$}, this then yields $\delta E_{I}=\int n_{I}({\bf r})\delta v_{\rm ext}({\bf r})d^{3}r$, where $n_{I}({\bf r})=\sum_{\sigma_{1}..\sigma_{N}}\int|\Psi_{I}({\bf r}\sigma_{1},{\bf r}_{2}\sigma_{2}...{\bf r}_{N}\sigma_{N})|^{2}d^{3}r_{2}..d^{3}r_{N}$ is the density of the unperturbed excited state $\Psi_{I}$. Thus, knowledge of the excited state energy and how it varies with the potential yields the excited-state density itself:
@@
Eq. 1 provides the cornerstone for obtaining excited state densities from TDDFT response: while ground-state DFT provides the ground-state energy $E_{0}$, the TDDFT linear response formalism provides excitation energies $\omega_{I}$, so taking the functional derivative with respect to the external potential of the sum of these yields the excited-state density, exact in principle and dependent on functional approximations made in practise, as we shall see shortly. We note that Refs. 24, 25 recently used Eq. 1 to find the exact excited state densities of an asymmetric Hubbard dimer from an exact state-dependent excited-state energy functional. Here instead we obtain the energy from TDDFT. Thus, we write the excited state energy in two parts:
@@
where $E_{0}=\left\langle\Psi_{0}|\hat{T}+\hat{W}|\Psi_{0}\right\rangle+\int drn_{0}(r)v_{\mathrm{ext}}(r)$ (with $\hat{T}$ and $\hat{W}$ as kinetic and electron-interaction operators, $\Psi_{0}$ the ground-state wavefunction, and $n_{0}$ the ground-state density), and $\omega_{I}=E_{I}-E_{0}$ is the excitation frequency of the excited state $I$. The latter are eigenvalues of the TDDFT matrix equation 26, 27, 28, 29:
@@
where the matrix elements of $\Omega(\omega)$ are expressed in the single-excitation basis, $q=i\rightarrow a,\;q^{\prime}=j\rightarrow b$, for spin-saturated systems as
@@
and $\Phi_{q}(\mathbf{x})=\phi_{i}(\mathbf{x})\phi_{a}(\mathbf{x})$ represents a KS single-excitation transition density. We note that while Eq. 4 represents a matrix equation in the single excitation basis, de-excitations are included due to the squared nature of the matrix 26, 29
Evaluating the derivative in Eq. 1 from Eq. 2 we have
@@
obtaining the derivative from Eq. 3 in the following way. Multiplying by $\mathbf{G}_{I}^{\dagger}$ on the left, and then taking the derivative, yields the left-hand side as
@@
$\Delta n_{I}({\bf r})=\frac{1}{2\omega_{I}}\frac{\delta\omega^{2}_{I}}{\delta v_{\text{ext}}({\bf r})}=\frac{1}{2\omega_{I}}\mathbf{G}_{I}^{\dagger}\frac{\delta\Omega(\omega_{I})}{\delta v_{\rm ext}({\bf r})}\mathbf{G}_{I}/{\mathcal{N}}$ (11)
where $\mathcal{N}=\mathbf{G}_{I}^{\dagger}\mathbf{G}_{I}$}. This expression gives us the exact density-difference, once we compute the functional derivative of the TDDFT matrix.
While Eq. 11 is independent of the choice of normalization $\mathcal{N}$} of the response eigenvectors $\mathbf{G}_{I}$ (as it must be from Eq. 8), we will consider the particular normalization derived in Ref. 26 which directly relates the eigenvectors to the oscillator strengths via $f_{I}=\frac{2}{3}\omega_{I}|\langle\Psi_{0}|{\bf r}|\Psi_{I}\rangle|^{2}=\frac{2}{3}\left|\sum_{q,q^{\prime}}\mathbf{r}_{q}^{\dagger}S_{q,q^{\prime}}^{-1/2}G_{I,q^{\prime}}\right|^{2}$ 27, where $S$ is the diagonal matrix $S_{qq^{\prime}}=\delta_{qq^{\prime}}\frac{1}{\nu_{q}}$. Ref. 26 showed that the required normalization condition is
@@
Solving this for $\mathcal{N}$} and inserting into Eq. 11 yields
$\Delta n_{I}({\bf r})=\frac{1}{2\omega_{I}}\left(\mathbbm{1}+\mathbf{G}_{I}^{\dagger}\left.\frac{\partial\Omega(\omega)}{\partial(\omega^{2})}\right|_{\omega=\omega_{I}}\mathbf{G}_{I}\right)^{-1}\mathbf{G}_{I}^{\dagger}\frac{\delta\Omega(\omega_{I})}{\delta v_{\text{ext}}({\bf r})}\mathbf{G}_{I}.$ (13)
We note that the functional derivative $\frac{\delta\Omega(\omega_{I})}{\delta v_{\rm ext}({\bf r})}$ has two contributions evident from Eq. 4: one is from the variation of KS orbitals and eigenvalues $\{\phi_{i}({\bf r}),\epsilon_{i}\}$ that explicitly appear in Eq. 4 and from the variation of the functional-argument $n_{0}$ of the $f_{\scriptscriptstyle\rm XC}$ kernel under small changes of the external potential, and the other is from the possible frequency-dependence of the xc kernel $f_{\scriptscriptstyle\rm XC}(\omega)$ because this is evaluated at the solution point of Eq. 3, $\omega_{I}$, which depends on $v_{\rm ext}({\bf r})$, $\frac{\delta\omega_{I}}{\delta v_{\rm ext}({\bf r})}$. Within the commonly-used adiabatic approximation, the TDDFT matrix is frequency-independent and so $\frac{\delta f_{\scriptscriptstyle\rm XC}(\omega)}{\delta\omega}=0$ and $\frac{\partial\Omega}{\partial(\omega^{2})}$ in Eq. 13 is zero. In the general case, frequency-dependence of the exact kernel brings non-linearity to Eq. 3, resulting therefore in more solutions than the size of the KS single-excitation matrix. This is essential to correctly get states corresponding to states of double-excitation character 30, 31, 32, 33, 34, which are not accessible with an adiabatic approximation. While the frequency-dependence of the $f_{\scriptscriptstyle\rm XC}$ kernel also affects the density-difference dependence through Eq. 13, this equation is, however, still linear, as the matrix is evaluated at the solution frequency $\omega_{I}$, yielding a unique and well-defined density-difference in Eq. 8.
A simplification arises once we separate the two contributions to the functional derivative of the $f_{\scriptscriptstyle\rm XC}$ kernel that were mentioned above:
@@
where the second term is only non-zero in the case of non-adiabatic functionals. Transferring this separation to the full $\Omega(\omega)$ matrix, and inserting into Eq. 13, we find
@@
There are two advantages of using Eq. 15. First, with the normalization Eq. 12 the derivative of Eq. 11 because the derivative on the right hand side of the resulting equation Eq. 15 no longer includes the variation of the solution point $\omega_{I}$ under $v_{\rm ext}$. Second, Eq. 15 lends a physical interpretation, in that the factor $|G_{I,q}|^{2}$ multiplying each diagonal term of the derivative matrix can be interpreted as a fraction of the single excitation $q$ in the density difference of the true excited state $I$. We will use Eq. 15 in what follows.
Eq. 15 (or Eq. 11) defines the exact excited-state densities once the terms of the derivative matrix $\frac{\delta\Omega_{qq^{\prime}}(\omega)}{\delta v_{\rm ext}({\bf r})}$ are calculated. When KS states are considered real, one finds:
@@
$+4\sqrt{\nu_{q}\nu_{q^{\prime}}}\left(\sum_{p\neq i}^{\infty}\frac{1}{\epsilon_{i}-\epsilon_{p}}{f_{\scriptscriptstyle\rm HXC}}_{,\,pa,\,jb}\left(\omega\right)\Phi_{ip}(\mathbf{x})+\sum_{p\neq a}^{\infty}\frac{1}{\epsilon_{a}-\epsilon_{p}}{f_{\scriptscriptstyle\rm HXC}}_{,\,ip,\,jb}\left(\omega\right)\Phi_{pa}(\mathbf{x})\right$.
$\left.\left.+\sum_{p\neq j}^{\infty}\frac{1}{\epsilon_{j}-\epsilon_{p}}{f_{\scriptscriptstyle\rm HXC}}_{,\,ia,\,pb}\left(\omega\right)\Phi_{jp}(\mathbf{x})+\sum_{p\neq b}^{\infty}\frac{1}{\epsilon_{b}-\epsilon_{p}}{f_{\scriptscriptstyle\rm HXC}}_{,\,ia,\,jp}\left(\omega\right)\Phi_{pb}(\mathbf{x})\right)\right\}\left(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}\right)^{-1}(\mathbf{x},{\bf r})$
@@
Section 1 in the Supplementary Material contains the details of the derivation of Eq. 16. Here we have used the following short-hands. First, we note that when $\chi,\chi_{\scriptscriptstyle\rm S},f_{\scriptscriptstyle\rm HXC}$ are written without any frequency-dependence, this indicates they are to be evaluated at the static limit $\omega=0$: $\chi({\bf r},{\bf r}^{\prime})=\frac{\delta n({\bf r})}{\delta v_{\rm ext}({\bf r}^{\prime})}|_{n=n_{0}},\chi_{\scriptscriptstyle\rm S}({\bf r},{\bf r}^{\prime})=\frac{\delta n({\bf r})}{\delta v_{\scriptscriptstyle\rm S}({\bf r}^{\prime})}|_{n=n_{0}},f_{\scriptscriptstyle\rm HXC}=f_{\scriptscriptstyle\rm HXC}[n_{0}]({\bf r},{\bf r}^{\prime})=\frac{\delta v_{\scriptscriptstyle\rm HXC}[n]({\bf r})}{\delta n({\bf r}^{\prime})}|_{n=n_{0}}$. In contrast, $f_{\scriptscriptstyle\rm HXC}(\omega)=f_{\scriptscriptstyle\rm HXC}[n_{0}]({\bf r},{\bf r}^{\prime},\omega)$ is defined as in Eq. 6. The KS approximation to the single-excitation density difference is denoted $\Delta n^{\scriptscriptstyle\rm KS}_{q}({\bf x})=\phi_{a}^{2}({\bf x})-\phi_{i}^{2}({\bf x})$, while $\Phi_{mn}({\bf x})=\phi_{m}({\bf x})\phi_{n}({\bf x})$ is defined in the same way as the the KS transition density earlier was, except now generalized since $\phi_{m},\phi_{n}$ could both be occupied or unoccupied orbitals. Note that $p$ in the sums run over all states, occupied and unoccupied, so this expression probes the xc kernel matrix elements (Eq. 5) in a much larger range than needed in the calculation of the excitation spectra; the latter involves matrix elements only between single excitations. The real-space matrix $(\mathbbm{1}-f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S})^{-1}({\bf x},{\bf r})$ is defined using the static $f_{\scriptscriptstyle\rm HXC}$ kernel and KS response function, and the product $f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}$ indicates the convolution, $f_{\scriptscriptstyle\rm HXC}\chi_{\scriptscriptstyle\rm S}({\bf x},{\bf r})=\int d^{3}x^{\prime}f_{\scriptscriptstyle\rm HXC}({\bf x},{\bf x}^{\prime})\chi_{\scriptscriptstyle\rm S}({\bf x}^{\prime},{\bf r})$. Finally, $\tilde{g}_{{\scriptscriptstyle\rm XC},qq^{\prime}}({\bf x})=\int d^{3}x^{\prime}d^{3}x^{\prime\prime}\tilde{g}_{\scriptscriptstyle\rm XC}[n_{0}](\mathbf{x}^{\prime},\mathbf{x}^{\prime\prime},\mathbf{x},\omega)\Phi_{q}({\bf x}^{\prime})\Phi_{q^{\prime}}({\bf x}^{\prime\prime})$ is the matrix in KS single-excitation space of
@@
Eqs. 1516, once added to the ground-state density, give in-principle the exact excited-state density for the state $I$. As such Eq. 16 must integrate to zero. To see that it does, consider the Taylor expansions of the inverse
@@
We note that Refs. 35, 36, 4, 5 begin instead with finding stationary points of a Lagrangian that expresses the variational property of the excited state energy, working in the KS single-excitation subspace to directly obtain matrix equations for the so-called $Z$-vector that contains relaxation contributions to a difference density matrix; solving these equations yields excited state gradients and other properties. Our Eqs. 1516 are equivalent but instead yields the densities directly in real-space, and moreover, applies also to the case of non-adiabatic kernels.